OpenStudy (caitlinnr14):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (caitlinnr14):

help please!

OpenStudy (anonymous):

\((n)+(n+1)+(n+2) = 267\) Solve for n

OpenStudy (caitlinnr14):

??????

OpenStudy (anonymous):

\((n)+(n+1)+(n+2) = 267 \\ \textrm{Combining like terms, we get } \\ 3n + 3 = 267\)

OpenStudy (caitlinnr14):

ohhhh thx @KevinOrr